# Undesirable Memorization in Large Language Models: A Survey

### v2

@@ IV-E Decoding methods
IV-F Fine-tuning and transfer learning
IV-G Training process dynamics
Zhang et al. [71], Kandpal et al. [23] show that memorization grows consistently with the number of training epochs, which makes sense, as more epochs push the model to potential overfitting. Jagielski et al. [93] demonstrate that examples seen during the earlier stages of training are less prone to memorization and rather they are forgotten over time. These findings indicate that memorization increases with more training, while early-seen examples being more likely to be forgotten. Leybzon and Kervadec [94] show higher memorization rates happens early and late in training, with lower rates mid-training. Therefore, suggesting that placing sensitive data in the middle stages of training could reduce its vulnerability to extraction attacks.
IV-H Forgetting mechanisms
Kirkpatrick et al. [100] initially introduced “catastrophic forgetting” in the context of neural networks and continual learning. They propose a method to protect important model weights to retain knowledge. This approach has been effective in maintaining performance on older tasks, even after long periods of non-use. Tirumala et al. [25] observe the forgetting mechanisms of a special batch through the learning process and show that it follows an exponential degradation, reaching a constant value baseline. They show that the mentioned baseline scales with the model size. Jagielski et al. [93] address the dual phenomena of memorization and forgetting in LLMs through stronger privacy attacks and several strategies for measuring the worst-case forgetting of the training examples. The study introduces a method to measure to what extent models forget specific training data, highlighting that standard image, speech, and language models do indeed forget examples over time, though non-convex models might retain data indefinitely in the worst case. The findings suggest that examples from early training phases, such as those used in pre-training large models, might enjoy privacy benefits but could disadvantage examples encountered later in training.
IV-I Explainability and interpretability
Huang et al. [101] study verbatim memorization in LLMs using controlled pre-training with injected sequences. They find that memorization requires significant repetition, increases in later checkpoints, and is tied to distributed model states and general language modeling capabilities. Their stress tests show unlearning methods often fail to remove memorized information without degrading model performance, highlighting the difficulty of isolating memorization. Haviv et al. [102] propose a framework to probe how memorized sequences are recalled in transformers, showing that memory recall follows a two-step process: early layers promote the correct token in the output distribution, while upper layers amplify confidence. They find that memorized information is primarily stored and retrieved in early layers. Similarly, Dankers and Titov [103] investigate where memorization occurs across model layers, demonstrating that memorization is a gradual and task-dependent process rather than localized to specific layers. Using centroid analysis, they show that deeper layers contribute more to memorization when models generalize well to new data. Stoehr et al. [104] show that memorization in LLMs, while distributed across layers, is driven by distinct gradients in lower layers and influenced by a low-layer attention head focusing on rare tokens. Perturbation analysis reveals that distinctive tokens in a prefix can corrupt entire continuations, and memorized sequences are harder to unlearn and more robust to corruption than non-memorized ones. Chen et al. [105] reveal interpretability insights into memorization by identifying clustering of sentences with different memorization scores in the embedding space and observing an inverse boundary effect in entropy distributions for memorized and unmemorized sequences. They also demonstrate that hidden states of LLMs can be used to predict unmemorized tokens, shedding light on the dynamics of memorization within the model. These findings offer insights into the mechanisms of memorization, enhancing interpretability and guiding future research.
@@ IV-I Explainability and interpretability
Model capacity Larger models memorize more [39, 25]
Training data Duplicated data amplifies memorization [77, 23, 25]
@@
Tokenization Bigger tokenizer vocabulary leads to more memorization [85]
@@
Training process Earlier phases of training are less prone to memorization [71, 23, 93]
Forgetting mechanisms Forgetting follows an exponentially decaying curve [25, 93]
TABLE II: Influencing factors and dynamics of memorization and their key findings (discussed in Section IV)
@@ V Mitigating memorization: strategies and techniques
As discussed in earlier sections, memorization is influenced by a range of factors and it is sometimes necessary for the learning process [111, 112]. However, in scenarios where memorization could lead to privacy concerns or security vulnerabilities, some methods could be employed to mitigate its impact. In these cases, specific strategies are utilized to limit the retention of sensitive information, ensuring that potential risks related to data exposure or misuse are minimized.
V-A Data de-duplication
Lee et al. [77] run exact matching and approximate matching (MiniHash) de-duplication algorithms on the C4 [113], RealNews [114], LM1B [115], and Wiki40B [116] datasets and show that these datasets contain up to 13.6% near duplicates and up to 19.4% exact duplicates.
@@ V-A Data de-duplication
V-B Differential privacy
@@
V-C Unlearning methods
@@
Unlearning methods offer effective strategies to mitigate memorization by selectively removing specific information from trained models without the need for full retraining. However, these methods are generally designed to target specific pieces of information, which means they rely on the ability to identify what the model has memorized beforehand. This reliance poses a challenge, as determining the exact content a model has memorized can be difficult and may limit the applicability of unlearning techniques for general mitigation against memorization.
V-D Heuristic Methods
Liu et al. [130] propose the alternating teaching method, a teacher-student framework where multiple teachers trained on disjoint datasets supervise a student model in an alternating fashion to reduce unintended memorization. This approach demonstrates superior privacy-preserving results on the LibriSpeech dataset [131] while maintaining minimal utility loss when sufficient training data is available. Ippolito et al. [26] propose “MemFree Decoding,” a novel sampling strategy to reduce memorization during text generation by avoiding the emission of token sequences matching the training data. They demonstrate its effectiveness across multiple models, including GPT-Neo [132] and Copilot [133], significantly lowering generation similarity with the training data. However, they also reveal that a simple style transfer in prompts can bypass this defense, highlighting its limitations. Similarly, Borec et al. [107] investigate the impact of nucleus sampling on memorization, showing that while increasing the nucleus size slightly reduces memorization, it only provides modest protection. They also highlight the distinction between “hard” memorization, involving verbatim reproduction, and “soft” memorization measured by the ROUGE similarity metric [134], where generated outputs echo the training data without exact replication. Hans et al. [135] introduce the “goldfish loss,” a subtle modification to the next-token training objective where randomly sampled subsets of tokens are excluded from the loss computation. This prevents models from memorizing complete token sequences, significantly reducing extractable memorization while maintaining downstream performance.
@@ VI Future directions and open challenges
VI-A Balancing performance and privacy in LLMs
As discussed earlier, memorization in LLMs can also have beneficial uses, such as improving factual recall and enhancing performance in specific tasks such as question answering. Privacy and copyright issues are some of the major concerns here [19, 20, 11]. Privacy-enhancing technologies such as multi-party computing, homomorphic encryption, and differential privacy, which can be employed to mitigate memorization risks and prevent sensitive data exposure, often come at the cost of reduced model performance, as they limit the model’s ability to retain precise information or perturb the model’s output. To fully harness the benefits of memorization while safeguarding against privacy breaches, future research needs to focus on strategies that balance these competing goals. This includes developing techniques that protect sensitive data and intellectual property without significantly degrading the model’s accuracy and utility. Such efforts will be key to ensuring both legal compliance and high-performance outcomes in LLMs.
@@ VI-D Memorization in specific contexts
Multilingual Large Language Models (xLM)
xLMs are trained to interpret and generate text in multiple languages. Trained on extensive datasets that include various languages, these models develop language-agnostic representations [142]. Despite their potential, xLMs often face challenges related to data scarcity for low-resource languages, leading to performance disparities [143]. This necessitates ongoing research to enhance their efficacy and fairness across all languages. In line with the significant effect of training data on memorization (discussed in section IV-B), for future work, we propose to investigate whether a low-resource language setting is more prone to memorization by comparing it to English scenarios. Addressing this research direction is essential not only for understanding the privacy implications associated with LLMs and data leakage in low-resource settings but also for ensuring AI safety in societies using LLMs with languages that have fewer resources.

### v3

@@ IV-C2 Decoding methods
Model capacity Larger models memorize more. [17, 140]
Training data Duplicated data amplifies memorization. [77, 66, 140]
@@
Tokenization Bigger tokenizer vocabulary leads to more memorization. [70]
@@
Training process Earlier phases of training are less prone to memorization. [162, 66, 64]
Forgetting mechanisms Forgetting follows an exponentially decaying curve. [140, 64]
TABLE II: Factors of influence and dynamics of memorization and their key findings (discussed in Section IV)
@@ V Mitigating memorization: strategies and techniques
As discussed in earlier sections, memorization is influenced by a range of factors and it is sometimes necessary for the learning process [37, 15]. However, in scenarios where memorization could lead to privacy concerns or security vulnerabilities, some methods could be employed to mitigate its impact. In these cases, specific strategies are utilized to limit this phenomenon, ensuring that potential risks related to data exposure or misuse are minimized. In this section, we provide a holistic overview of such methods, categorized in three groups: data-level, training-time, and post-training approaches.
V-A Data-level approaches
As discussed in section IV-B1, duplicated training data is one of the major factors of memorization. Lee et al. [77] run exact matching and approximate matching (MiniHash) de-duplication algorithms on the C4 [31], RealNews [159], LM1B [23], and Wiki40B [44] datasets and show that these datasets contain up to 13.6% near duplicates and up to 19.4% exact duplicates.
@@ V-A Data-level approaches
V-B Training-time approaches
V-B1 Differential privacy
@@
V-B2 Promoting Reasoning
A growing line of work aims to disentangle and control the interplay between reasoning and memorization in LLMs, offering tools to reduce memorization-driven failures and benchmark inflation. Salido et al. [121] introduce “None of the Others (NOTO)”, a simple modification to multiple-choice evaluations where the correct answer is replaced by “None of the others,” forcing models to reason by elimination rather than rely on memorized answer patterns. The large accuracy drops observed under NOTO reveal strong dependence on memorization and contamination in standard benchmarks. Complementary mechanistic insights are provided by Hong et al. [54], who show that the shift between reasoning and memorization is mediated by a single linear direction in the model’s residual stream. By identifying and manipulating this reasoning feature, they demonstrate causal control over whether a model behaves like a reasoner or retrieves memorized content, suggesting targeted intervention as a potential mitigation path. Du et al. [32] find that memorization is not a distinct or isolated mechanism but emerges through the same pathways that support reasoning. Even when models memorize noisy labels, they still execute intermediate reasoning steps, indicating that memorization is distributed and intertwined with reasoning. This challenges assumptions that memorization can be cleanly isolated, and motivates mitigation strategies that focus on shaping reasoning pathways rather than simply suppressing memory lookup. Finally, Li and Goyal [80] propose Memory Conditioned Training (MCT) to integrate new knowledge into LLMs. While standard updates improve memorization, reasoning over updated facts remains difficult. MCT prepends “memory tokens” during training, helping models recall new facts while still performing reasoning, highlighting the challenge of aligning memorization and reasoning in dynamic knowledge settings.
Together, these studies highlight strategies that encourage reasoning while managing memorization, providing a multi-faceted approach to mitigating failures driven by memorization.
V-B3 Training intervention methods
Liu et al. [85] propose the alternating teaching method, a teacher-student framework where multiple teachers trained on disjoint datasets supervise a student model in an alternating fashion to reduce unintended memorization. This approach demonstrates superior privacy-preserving results on the LibriSpeech dataset [106] while maintaining minimal utility loss when sufficient training data is available. Hans et al. [46] introduce the “goldfish loss,” a subtle modification to the next-token training objective where randomly sampled subsets of tokens are excluded from the loss computation. This prevents models from memorizing complete token sequences, significantly reducing extractable memorization while maintaining downstream performance. Ghosal et al. [41] propose “Memorization Sinks,” a training-based method that isolates memorized content into dedicated neurons, preventing it from interfering with general language knowledge and reducing the risk of memorizing sensitive information.
V-C Post-training approaches
V-C1 Unlearning methods
@@
V-C2 Model editing
Ruzzetti et al. [120] propose “Private Memorization Editing (PME),” a technique that detects memorized PII in a trained LLM and edits the model’s internal memory to remove that PII, thereby reducing the risk of privacy leakage while preserving the model’s general language capabilities. In a similar line of thought, Ni et al. [101] introduce a weight‑pruning based method that provides a unified framework to tune a model’s memorization: by gradient‑guided pruning, one can suppress or amplify memorization depending on privacy vs utility needs, while balancing general language performance.
V-C3 Decoding-based methods
Ippolito et al. [61] propose “MemFree Decoding,” a novel sampling strategy to reduce memorization during text generation by avoiding the emission of token sequences matching the training data. They demonstrate its effectiveness across multiple models, including GPT-Neo [10] and Copilot [42], significantly lowering generation similarity with the training data. However, they also reveal that a simple style transfer in prompts can bypass this defense, highlighting its limitations. Similarly, Borec et al. [13] investigate the impact of nucleus sampling on memorization, showing that while increasing the nucleus size slightly reduces memorization, it only provides modest protection. They also highlight the distinction between “hard” memorization, involving verbatim reproduction, and “soft” memorization measured by the ROUGE similarity metric [83], where generated outputs echo the training data without exact replication.
Remark.
Data-level mitigations can reduce memorization but have limitations: deduplication may hurt task performance, and data sanitization is not fully reliable [87, 113]. Training interventions, including DP, alternating teaching, and other heuristic approaches, offer stronger control during training. DP provides formal privacy guarantees but depends on hyperparameters and can still leave residual memorization. Other methods may interfere with long-range dependencies or internal representations, affecting downstream reasoning.
Post-training methods, such as unlearning, model editing, and decoding-based strategies, allow targeted mitigation without retraining, but they rely on identifying memorized content, risk unintended side effects, and cannot fully address hidden memorization or reasoning influences.
Together, these approaches highlight that while mitigation is possible, no single method fully eliminates memorization, and careful combination is needed to balance privacy, reasoning, and performance.
@@ VI Future directions and open challenges
VI-A Balancing memorization benefits and privacy risks
As discussed earlier, memorization in LLMs can have beneficial uses, such as improving factual recall and enhancing performance in specific tasks like question answering. At the same time, privacy and copyright concerns remain major challenges [99, 134, 156]. Interestingly, recent work challenges the belief that mere verbatim memorization leads to privacy leakage; concluding that what drives sensitive information exposure is how memorized content is utilized or revealed by the model [122]. On the other hand, standard privacy-enhancing technologies, such as multi-party computation, homomorphic encryption, and differential privacy, can mitigate memorization risks, but often reduce model performance by limiting precise retention or perturbing outputs. Future research should focus on strategies that balance memorization benefits with privacy protection, developing techniques that safeguard sensitive data and intellectual property without significantly degrading accuracy or utility. Such approaches are essential for ensuring both high-performance outcomes and legal compliance in LLM deployment.
@@ VI-D Memorization in specific contexts
Multilingual Large Language Models (xLLM)
xLLMs are trained to interpret and generate text across multiple languages, often relying on large multilingual corpora to develop language-agnostic representations [115]. Yet, they commonly suffer from data scarcity in low-resource languages, resulting in notable performance disparities [155]. Recent work highlights that memorization in multilingual models is not uniform across languages. Luo et al. [89] show that memorization patterns are shaped by cross-lingual similarities rather than data volume alone, and Satvaty et al. [124] demonstrate that lower-resource languages show different memorization patterns when compared to the English language. These findings motivate future work to more systematically analyze how memorization and privacy risks differ across languages and to uncover the underlying factors driving these differences, and to develop methods that effectively mitigate them.